OpenStudy (anonymous):

what is f(g(2) if f(x)= 2x+3, and g(x)= 3x+14

OpenStudy (anonymous):

43

OpenStudy (anonymous):

can you explain it and show work

OpenStudy (anonymous):

g(2)= 3(2) + 14=20 f(20)= 2(20) +3=43